# Machine learning classification of black holes in the mass–spin diagram

### v1

@@ II Black hole mass-spin diagrams
We note that one can construct analogous diagrams, highlighting the ultimate flexibility of this framework, to adapt it to specific use cases. For example, one could transform the spin magnitude axis into a redshift axis given a cosmological model for the spin redshift-dependence for future gravitational wave detectors. This diagram can also be extended to three dimensions by including a new axis or colormap for the spin orientation if the black holes exist in binaries, or relevant astrophysical parameters as in Fig. 4. Other quantities that depend on the black hole mass(es) and spin(s), such as the recoil kick velocity in black hole formation from gravitational collapse or in black hole mergers from binary systems, would appear as contours on the mass-spin diagram. Even though the sensitivity of the LVK detectors vanishes beyond $z\approx 0.1$, the masses and spins and formation rates of stellar-mass black hole binaries depend on the redshift and metallicity dependent star formation rate evolution [105, 106]. The black hole spins are long expected to be key tracers of that evolution and to act as signposts for certain formation pathways and channels, and can thus be used as a proxy -or an intermediate- for the redshift in some cases such as strongly accreting populations, allowing for new comparisons between X-ray binaries and merging binaries [107, 108]. Future GW detectors will open the horizon to much higher distances, ie $z\lesssim 10$, indicating a need for new theoretical tools that encapsulate the wealth of scientific information these detectors will provide. The mass-spin diagram offers a natural framework to study the combined EM and GW black hole populations, as they generally originate in diverse environments and at varying redshifts, helping to empower future multi-messenger astrophysics studies of black holes of varying masses [109]. Indeed, this multi-messenger concept can be extended to non-binary sources of GWs such as continuous wave emission from Galactic isolated neutron stars with LVK or ultra-compact binaries with LISA.
Figure 5: A simple example of a black hole mass gap, i.e. no intermediate-mass black holes, and distinct and separated mass–spin correlations for stellar-mass and supermassive black holes. Given to a bare unsupervised clustering algorithm [40], it may or may not find the two distinct classes of black holes shown by the red and blue circles. In this case, k-means (right panel) does not accurately reproduce the two classes but a Gaussian mixture model (left panel) clusters with 100% accuracy.
Figure 6: Unsupervised spectral clustering [40] on the natal black hole mass–spin correlations from Fig. 3. The stellar-mass, intermediate-mass, and supermassive regimes are shown by blue circles, red pluses, and green x’s, respectively. Black vertical lines show the two mass boundaries imposed in the true dataset, which the algorithm struggles to correctly reproduce.
@@ III.1 The Black Hole Continuum
Intermediate-mass black holes (IMBHs; $\approx 10^{2}\,\rm\,M_{\odot}$ to $10^{5}\,\rm\,M_{\odot}$) occupy a contentious region between well-established stellar-mass and supermassive regimes. EM candidates in ultra-luminous and hyper-luminous X-ray sources, dynamical globular clusters, or dwarf galaxies are numerous but have yet to be confirmed confidently [Greene2020]. Meanwhile, recent GW events with LVK (e.g., GW190521, GW200220, GW231123) concretely confirm the presence of lower-end IMBHs, offering glimpses into this apparent mass gap.
@@
First, we generate a toy dataset comprising two populations: stellar-mass black holes with spin magnitudes uniformly distributed between 0 and 1, and supermassive black holes with spins uniformly between 0.9 and 1. Figure 5 shows how a Gaussian mixture model on this bimodal distribution successfully separates stellar and supermassive populations based solely on mass-spin correlations, whereas the bare k-means struggles. This hints toward a potential for machine learning methods to reproduce—and ultimately refine—existing classifications, anticipate new categories as datasets grow, and identify selection biases inherent in human-based taxonomy.
Building on this preliminary test, next we construct a more astrophysically realistic black hole continuum in Figure 6. Here, we sample three initial mass functions (the same ones used to population Fig. 3) corresponding to stellar, intermediate-mass (uniform), and supermassive populations, and we assign each regime spin magnitude distributions of increasing concentration: uniform in [0,1] for stellar, bimodal for intermediate, and high spins ($\chi\gtrsim 0.8$) for supermassive black holes. This theoretical dataset composed of three mass-spin regimes challenges unsupervised learning algorithms to discover three clusters without knowledge of their labels.
Gaussian mixture and k-means models fail to recover the three classes reliably as they both cluster the stellar- and intermediate-mass black holes into a single class. In contrast, spectral clustering (with nearest neighbor affinity) operates on the full mass–spin manifold and accurately recovers the tripartite division, as shown by the three color-marker pairs in Fig. 6 where the vertical black lines indicate the true mass boundaries assumed in constructing the theoretical dataset. The misclassification of edge cases reflects the spin distinctions we imposed, but more importantly it demonstrates that such algorithms can detect latent structure in complex data which may arise from selection effects or data analysis assumptions. Also, it provides further motivation for the development of a black hole continuum mass function, which is an active area of research [110] and would present an interesting challenge for these strategies. We note that this classification was performed without latent space embedding prior to clustering, implying that an analysis which includes latent space transformation may perform even better, which we leave to future work.
@@ III.2 Deep Learning of Stellar-mass Black Holes
Instead of the output of simulations, for case 3 we construct distributions of mass and spin to simulate two theoretical correlations in the $\chi_{\rm eff}$$q$ parameter space corresponding to isolated binaries or dynamically formed binaries. First we sample the mass ratio $q$ uniformly in $(0.2,0.9998)$ and total mass $M$ uniformly in $(30,100)\,M_{\odot}$, from which the component masses $m_{1}$ and $m_{2}$ are derived. For the dynamical-cluster binaries, spin magnitudes $\chi_{1}$ and $\chi_{2}$ are set to be relatively high ($\sim 0.7$) and decreasing linearly with $q$, and spin-orbit misalignment cosines $\cos\theta$ are drawn uniformly from [1,1][-1,1]. The effective spin $\chi_{\mathrm{eff}}$ is computed from these quantities. For the isolated binaries, $\chi_{\mathrm{eff}}$ is computed from spin magnitudes that are lower on average and given by $\mathcal{N}(0.5,0.1)$, and spin-orbit misalignments $\cos\theta$ that are mostly aligned and modeled as decreasing linearly with $q$ and with spread given by $\mathcal{N}(0.5,0.1)$. In this case, samples are stored with population labels and written to a JSON file. Another main difference for case 3 is that the latent space is defined by $\chi_{\rm eff}$$q$ instead of the individual masses and spins, incorporating more information than using the binary masses and spins.
Using PyTorch [39] we implement and benchmark a neural network architecture which imposes a regularized latent structure, and we perform clustering with scikit-learn [40]. We do this in a modular, end-to-end machine learning framework to extract latent structure in simulated black hole populations and perform classification  [^4] All codes used in this study will be made publicly available upon publication. The workflow follows a standard extract–transform–load pipeline, where raw simulation outputs are extracted, processed, embedded into a compressed latent space with PyTorch neural net architecture, and clustered with the scikit-learn machine learning suite. We use an unsupervised classification scheme via a variational autoencoder (VAE) and K-means clustering, extending it to a semi-supervised classification framework, and a random forest supervised classifier. Raw data are split into 60% training, 20% test, and 20% validation sets containing black hole masses, spin magnitudes, and true class labels. As the performance is the same to within 1% between the training, test, and validation datasets, we only show the training results for brevity. All three datasets contain 100,000 black holes in each class, but results in Table 1 are similar for 10,000 black holes each.
@@
Table 1: Performance metrics of clustering with unsupervised classifier (k-means) and a supervised classifier (random forest) on the latent space representation from a variational autoencoder (VAE) for distinguishing the mass–spin correlations of the three black hole binary datasets described in the text. The metrics utilized here are the accuracy (i.e. the fraction of correctly labeled black holes), the adjusted Rand index (ARI), normalized mutual information (NMI), and F1-score.
@@
The classification and benchmarking results for the three datasets are shown in Table 1, where we only report the training results for brevity as the test and validation results are similar (i.e. $\approx 10$% difference). The results are similar for each of the three theoretical datasets. The unsupervised k-means clustering is unable to correctly learn the highly overlapping mass–spin distributions. This is demonstrated by the $\approx 0$ ARI and NMI indicating ineffective latent space learning. The consistent F1-score of $\sim 0.5$ for k-means across different hyperparameter settings combined with the near-0 ARI and NMI indicates that the k-means approach ends up with a random $\sim 50/50$ clustering assignment. We also confirm independently that the ARI and NMI are 0 ($\sim 0.3$) when unsupervised (supervised) clustering is performed directly on the raw input data rather than on the latent representation space, implying that the latent transformation is not redundant even with the same dimensionality as the raw data. We performed this analysis with a Gaussian mixture model instead of k-means and it similarly struggled to learn the latent parameter space. We explored this result by running a basic hyperparameter-space exploration in a gridded search to find an optimal solution, including exploring the latent space dimensionality (e.g., 4, 8, 16, 32).
Figure 7 demonstrates the difference in clustering in the trained VAE latent space between the unsupervised (top row panels) and supervised (bottom row panels) clustering of dataset 1., where the true and predicted data correspond to the left and right columns, respectively. Blue circles and orange squares correspond to the two spin models of Bavera et al. (2020) and Steinle & Kesden (2021), respectively. This visually confirms the challenge of latent space learning for the unsupervised approach. The success of the supervised random forest classifier (i.e.,  98% accuracy) is shown in the bottom row of Fig. 7 as the two predicted clusters are closely reproduced in the right panel with respect to the true clusters in the left panel. The dataset 1. has more distinct classes compared to datasets 2. and 3. And the spins of the two classes in dataset 1. correspond to the same black hole masses but in datasets 2. and 3. the spins of the two classes have different corresponding masses; both of these features contribute to the random forest performing better on dataset 1.
@@
While the unsupervised approach will require further investigation in future work, the success of the supervised random forest classifier indicates that a semi-supervised approach might show intermediate performance. Thus, we expanded the k-means pipeline for semi-supervised classification in which we assume some fraction has known labels and the training for the unlabeled fraction is the same as in the unsupervised pipeline but training for the labeled fraction employs an additional penalty in the loss function where points farther away from each other in the parameter space have less weight via cross-entropy loss in PyTorch. The result of the semi-supervised approach is sensitive to the initialization of random states in several components of the pipeline. A fuller analysis is beyond the scope of the present study, so we ran the semi-supervised k-means pipeline with many random initializations and a labeled fraction of 0.5 (chosen at random). The resultant best semi-supervised clustering had accuracy $\approx$ 0.6, f1-score $\approx$ 0.58, and NMI and ARI $\approx$ 0.45. Comparing with the fully unsupervised case, this demonstrates that the low-dimensional latent space transformation can improve clusterability. This indicates that higher dimensional mass–spin datasets could experience similar improvements in clustering by latent space representation learning such as utilized here. A path forward might be to utilize expert-informed analysis where a fraction of outliers in a dataset can be confidently labeled to guide training.
In the semi-supervised case, the class prediction depends on whether it is trained on $\chi_{\rm eff}$ and $q$ or on the masses and spins (i.e., a 2D physical space with a 2D latent space vs. a 4D physical space with a 2D latent space), and that these assumptions are important for cleanly learning the latent structure of observed binaries. The use of uncertainty in the event parameters is a key area of future development, which will benefit from other algorithms such as Bayesian Neural Networks that can quantify uncertainty and adaptively adjust relative importance of neurons. Such an analysis can be paired with hierarchical inference output [29], allowing the ML algorithms to operate directly on the posteriors rather than summary statistics (i.e. medians), enabling better utilization of astrophysical information for latent space learning, an effort we leave to future work. Our results indicate that combining the mass–spin diagram with machine learning contains potential for many applications in studying astrophysical black hole populations and distinguishing formation channels which is a great open problem in gravitational-wave astronomy, see e.g. [113, 114, 115, 116, 54, 117].
This motivates development of new unsupervised and semi-supervised learning methods, and fine-tuning of the approaches attempted here, for challenging astrophysical datasets from future GW and EM detectors, which we leave to a followup study [118]. In future work we will extend this framework to include higher dimensional input data, e.g. by adding the spin-orbit misalignments, binary mass ratio, and others to the black hole masses and spins, to develop new machine learning approaches to classifying astrophysical black holes in the mass–spin diagram. Another avenue of future work will use more realistic combined populations of black-hole binary formation channels in a cosmological setting, i.e., the redshift evolution across the mass–spin diagram.
@@ IV Summary
We present a new framework for visualizing and interpreting black hole populations with the mass–spin ($m-\chi$) diagram, analogous to the classic HR diagram in stellar astrophysics. While the HR diagram has long served as a cornerstone in our understanding of stellar evolution, a similarly intuitive and physically motivated diagnostic tool for dark compact objects can be found in the $m-\chi$ diagram. Capturing imprints of mass and angular momentum evolution, one can examine the time or redshift evolution and compare formation possibilities.
@@
First we examined the $m-\chi$ diagram in the context of the three black hole mass regimes: stellar-mass black holes formed from gravitational stellar core collapse, intermediate-mass black holes possibly formed through runaway collisions or direct collapse, and supermassive black holes from light and heavy seeds. We approximate redshift evolution using physically motivated mass and spin prescriptions, incorporating both super-Eddington accretion and stochastic processes. These black hole trajectories serve as tracks in the $m-\chi$ space, shaped by their accretion history and seed properties. For instance, stellar-mass BHs can spin up efficiently via fallback or disk accretion, while intermediate-mass black holes undergo diverse accretion and merger episodes, and supermassive black holes experience coherent spin-up in the early Universe and hierarchical galaxy mergers. The mass–spin trajectories carry a fossil record of evolutionary mechanisms, analogous to stellar tracks and clusters in the HR diagram, motivating inference of a black hole ’main sequence’ driven by accretion through cosmic evolution.
@@
Additionally, we provide a zoomed-in view on the stellar-mass black hole regime in a $m-\chi$ diagram constructed with data from the COMPAS binary evolution code. We post-process the black hole to obtain their spins with three models with subtle differences regarding the evolution of Wolf-Rayet tidal evolution in progenitor binaries of black hole mergers. The $m-\chi$ diagram exposes these differences in modeling assumptions motivating development of comprehensive and consistent spin evolution models. This will be most useful in the era of future terrestrial GW detectors, where catalogs of black hole binaries are expected to expand by at least 100 times in size resulting in tens of thousands of binaries [119, 120], consistent with the size of the populations utilized here.
We apply a suite of machine learning tools for unsupervised and supervised classification to a series of realistic and progressively challenging datasets. This demonstrates the utility of the mass–spin diagram for precisely classifying and interpreting black hole populations by their physical masses and spins. First we show how simple unsupervised algorithms can recover features in the mass–spin space, such as the intermediate-mass gap with k+means clustering and commonly used mass boundaries in a mass continuum with spectral clustering. Without latent space representations, these examples reveal the rich learning space available for these clustering algorithms to explore.
@@
Moving beyond simple clustering, we employ deep learning with a variational autoencoders (VAE), a type of neural network, to learn low-dimensional latent embeddings of three synthetic datasets for black hole masses and spins. After training the VAEs on standardized features using PyTorch, we extract the 2D latent spaces and apply an unsupervised algorithm, k-means clustering, and a supervised algorithm, random forest, for machine learning classification. In all three cases, the k-means approach does not learn the latent space representation and fails to accurately cluster the black holes according to their mass–spin correlations. This is clearly shown in Fig. 7 where the unsupervised latent space learning is shown in the top row panels compared to the supervised learning in the bottom panels for true (left column panels) vs predicted (right column panels). The random forest performs best for the first dataset, which is composed of two classes of black hole spins from models with differing assumptions about Wolf-Rayet tidal evolution, with near perfect accuracy and high ARI/NMI. This is because the two classes exhibit more distinct correlations in the mass–spin space compared to the other two datasets which have highly overlapping classes. In the case of dataset 2, this is simply the outcome of comparing raw simulation output from two state-of-the art models of black-hole binary formation channels; in the case of dataset 3, this is by construction to approximate an observed correlation in LVK data.
@@
These results showcase the utility of the mass–spin diagram for quantifying differences between spin evolution models and distinguishing formation channels of black-hole binaries. Our application of machine learning classification to the diagram reveals main challenges that lay ahead, motivating the development of advanced techniques to learn highly overlapping latent space representations and to include astrophysically informed priors or other information. Future work will identify how the motion of black holes across the mass–spin diagram can aid in their classification.

### v2

@@ II Black hole mass-spin diagrams
We note that one can construct analogous diagrams, highlighting the ultimate flexibility of this framework, to adapt it to specific use cases. For example, one could transform the spin magnitude axis into a redshift axis given a cosmological model for the spin redshift-dependence for future gravitational-wave detectors. This diagram can also be extended to three dimensions by including a new axis or colormap for the spin orientation if the black holes exist in binaries, or relevant astrophysical parameters as in Fig. 4. Other quantities that depend on the black hole mass(es) and spin(s), such as the recoil kick velocity in black hole formation from gravitational collapse or in black hole mergers of binary systems, would appear as contours on the mass-spin diagram. Even though the sensitivity of the LVK detectors vanishes beyond $z\approx 0.1$, the masses and spins and formation rates of stellar-mass black hole binaries depend on the redshift and metallicity dependent star formation rate evolution [106, 107]. The black hole spins are long expected to be key tracers of that evolution and to act as signposts for certain formation pathways and channels, and can thus be used as a proxy -or an intermediate- for the redshift in some cases such as strongly accreting populations, allowing for new comparisons between X-ray binaries and merging binaries [108, 109]. Future GW detectors will open the horizon to much higher distances, ie $z\lesssim 10$, indicating a need for new theoretical tools that encapsulate the wealth of scientific information these detectors will provide. The mass-spin diagram offers a natural framework to study the combined EM and GW black hole populations, as they generally originate in diverse environments and at varying redshifts, helping to empower future multi-messenger astrophysics studies of black holes of varying masses [110]. Indeed, this multi-messenger concept can be extended to non-binary sources of GWs such as continuous wave emission from Galactic isolated neutron stars with LVK or ultra-compact binaries with LISA.
Figure 5: A simple example of a black hole mass gap, i.e. no intermediate-mass black holes, and distinct and separated mass–spin correlations for stellar-mass and supermassive black holes. Given to bare unsupervised clustering algorithms [41], k-means (right panel) and a Gaussian mixture model (left panel) cluster with 100% accuracy.
Figure 6: Unsupervised spectral clustering [41] on the natal black hole mass–spin correlations from Fig. 3. The resultant clusters are shown by blue circles, red pluses, and green x’s. Black vertical lines show the two mass boundaries imposed in the true dataset between the stellar-mass, intermediate-mass, and supermassive regimes.
@@ III.1 The Black Hole Continuum
Intermediate-mass black holes (IMBHs; $\approx 10^{2}\,\rm\,M_{\odot}$ to $10^{5}\,\rm\,M_{\odot}$) occupy a contentious region between well-established stellar-mass and supermassive regimes. EM candidates in ultra-luminous and hyper-luminous X-ray sources, dynamical globular clusters, or dwarf galaxies are numerous but have yet to be confirmed confidently [21]. Meanwhile, recent GW events with LVK (e.g., GW190521, GW200220, GW231123) concretely confirm the presence of lower-end IMBHs, offering glimpses into this apparent mass gap.
@@
First, we generate a toy dataset comprising two populations: stellar-mass black holes with spin magnitudes uniformly distributed between 0 and 1, and supermassive black holes with spins uniformly between 0.9 and 1. Figure 5 shows how a Gaussian mixture model (left panel) and k-means (right panel) for clustering on this bimodal distribution successfully separates stellar and supermassive populations due to the simple mass-spin correlations. We perform this clustering in log-mass space to ensure proper feature scaling. This hints toward a potential for machine learning methods to reproduce—and ultimately refine—existing classifications, anticipate new categories as datasets grow, and identify selection biases inherent in human-based taxonomy.
Building on this preliminary test, next we construct a more astrophysically realistic black hole continuum in Figure 6. Here, we sample three initial mass functions (the same ones used to population Fig. 3) corresponding to stellar, intermediate-mass (uniform), and supermassive populations, and we assign each regime spin magnitude distributions of increasing concentration: uniform in [0,1] for stellar, bimodal for intermediate, and high spins ($\chi\gtrsim 0.8$) for supermassive black holes. This theoretical dataset can challenge unsupervised learning algorithms to discover three clusters without knowledge of their labels.
Spectral clustering (with nearest neighbor affinity), which operates on the full mass–spin manifold, can struggle to accurately recover the tripartite division, as shown by the three color-marker pairs in Fig. 6 where the vertical black lines indicate the true mass boundaries assumed in constructing the theoretical dataset. The misclassification of edge cases reflects the spin distinctions we imposed, but more importantly it demonstrates that such algorithms can detect latent structure in complex data which may arise from selection effects or data analysis assumptions. Also, it provides further motivation for the development of a black hole continuum mass function, which is an active area of research [111] and would present an interesting challenge for these strategies. We note that this classification was performed without latent space embedding prior to clustering, implying that an analysis which includes latent space transformation may perform even better, which we leave to future work.
@@ III.2 Deep Learning of Stellar-mass Black Holes
Instead of the output of simulations, for case 3 we construct distributions of mass and spin to simulate two theoretical correlations in the $\chi_{\rm eff}$$q$ parameter space corresponding to isolated binaries or dynamically formed binaries. First we sample the mass ratio $q$ uniformly in $(0.2,0.9998)$ and total mass $M$ uniformly in $(30,100)\,M_{\odot}$, from which the component masses $m_{1}$ and $m_{2}$ are derived. For the dynamical-cluster binaries, spin magnitudes $\chi_{1}$ and $\chi_{2}$ are set to be relatively high ($\sim 0.7$) and decreasing linearly with $q$, and spin-orbit misalignment cosines $\cos\theta$ are drawn uniformly from [1,1][-1,1]. The effective spin $\chi_{\mathrm{eff}}$ is computed from these quantities. For the isolated binaries, $\chi_{\mathrm{eff}}$ is computed from spin magnitudes that are lower on average and given by $\mathcal{N}(0.5,0.1)$, and spin-orbit misalignments $\cos\theta$ that are mostly aligned and modeled as decreasing linearly with $q$ and with spread given by $\mathcal{N}(0.5,0.1)$. Another main difference for case 3 is that the latent space is defined by $\chi_{\rm eff}$$q$ instead of the individual masses and spins, incorporating more information than using the binary masses and spins.
Using PyTorch [40] we implement and benchmark a neural network architecture which imposes a regularized latent structure, and we perform clustering with scikit-learn [41]. We do this in a modular, end-to-end machine learning framework to extract latent structure in simulated black hole populations and perform classification  [^4] Codes used in this study will be made publicly available upon publication. The workflow follows a standard pipeline, where raw simulation outputs are extracted, processed, embedded into a compressed latent space with PyTorch neural net architecture, and clustered with the scikit-learn machine learning suite. We use an unsupervised classification scheme via a variational autoencoder (VAE) and k-means clustering, extending it to a semi-supervised classification framework, and a random forest supervised classifier. Raw data are split into 60% training, 20% test, and 20% validation sets containing black hole masses, spin magnitudes, and true class labels. As the performance is the same to within 1% between the training, test, and validation datasets, we only show the training results for brevity. All three datasets contain 100,000 black holes in each class, but results in Table 1 are similar for 10,000 black holes each.
@@
Table 1: Performance metrics of clustering with an unsupervised classifier (k-means) and a supervised classifier (random forest) on the latent space representation from a variational autoencoder (VAE) for distinguishing the mass–spin correlations of the three black hole binary datasets described in the text. The metrics utilized here are the accuracy (i.e. the fraction of correctly labeled black holes), the adjusted Rand index (ARI), normalized mutual information (NMI), and F1-score.
@@
The classification and benchmarking results for the three datasets are shown in Table 1, where we only report the training results for brevity as the test and validation results are similar (i.e. $\lesssim 10$% difference). The unsupervised k-means clustering is unable to correctly learn the highly overlapping mass–spin distributions. This is demonstrated by the $\approx 0$ ARI and NMI indicating ineffective latent space learning. The consistent F1-score of $\sim 0.5$ for k-means across different hyperparameter settings combined with the near-0 ARI and NMI indicates that the k-means approach ends up with a random $\sim 50/50$ clustering assignment. We also confirm independently that the ARI and NMI are 0 ($\sim 0.3$) when unsupervised (supervised) clustering is performed directly on the raw input data rather than on the latent representation space. We also performed this analysis with a Gaussian mixture model instead of k-means and it performed similarly.
Figure 7 demonstrates the difference in clustering in the trained VAE latent space between the unsupervised (top row panels) and supervised (bottom row panels) clustering of dataset 1., where the true and predicted data correspond to the left and right columns, respectively. Blue circles and orange squares correspond to the two spin models of Bavera et al. (2020) and Steinle & Kesden (2021), respectively. This visually confirms the challenge of latent space learning for the unsupervised approach. The success of the supervised random forest classifier (i.e.,  98% accuracy) is shown in the bottom-row panels of Fig. 7, as the two predicted clusters are closely reproduced in the right panel with respect to the true clusters in the left panel. The dataset 1. has more distinct classes compared to datasets 2. and 3., and the spins of the two classes in dataset 1. correspond to the same black hole masses but in datasets 2. and 3. the spins of the two classes have different corresponding masses; both of these features contribute to the random forest performing better on dataset 1.
@@
While the unsupervised approach will require further investigation in future work, the success of the supervised random forest classifier indicates that a semi-supervised approach might show intermediate performance. Thus, we expanded the k-means pipeline for semi-supervised classification in which we assume some fraction has known labels and the training for the unlabeled fraction is the same as in the unsupervised pipeline but training for the labeled fraction employs an additional penalty in the loss function where points farther away from each other in the parameter space have less weight via cross-entropy loss in PyTorch. The result of the semi-supervised approach is sensitive to the initialization of random states in several components of the pipeline. , so we ran the semi-supervised k-means pipeline with 5 different initializations and a labeled fraction of 0.5. The resultant best semi-supervised clustering had accuracy $\approx$ 0.6, f1-score $\approx$ 0.58, and NMI and ARI $\approx$ 0.45. Comparing with the fully unsupervised case, this demonstrates that the low-dimensional latent space transformation can improve clusterability, but a full analysis is beyond the scope of the present study.
The use of uncertainty in the event parameters is a key area of future development, suggesting the use of methods like Bayesian Neural Networks that can quantify uncertainty in training and prediction. Such an analysis can be paired with hierarchical inference output [29], allowing the ML algorithms to operate directly on the posteriors rather than summary statistics (i.e. medians), enabling better utilization of astrophysical information for latent space learning. Our results indicate that combining the mass–spin diagram with machine learning contains potential for many applications in studying astrophysical black hole populations where distinguishing formation channels which is a great open problem in gravitational-wave astronomy, see e.g.’s [114, 115, 116, 117, 55, 118, 119]. This motivates development of new unsupervised and semi-supervised learning methods, and fine-tuning of the approaches attempted here, for challenging astrophysical datasets from future GW and EM detectors, which we leave to a followup study [120]. Another avenue of future work will use more realistic combined populations of black-hole binary formation channels in a cosmological setting, i.e., the redshift evolution across the mass–spin diagram.
@@ IV Summary
We present a new framework for visualizing and interpreting black hole populations with the mass–spin diagram, analogous to the classic HR diagram in stellar astrophysics. While the HR diagram has long served as a cornerstone in our understanding of stellar evolution, a similarly intuitive and physically motivated diagnostic tool for black holes can be found in the mass–spin diagram. Capturing imprints of mass and angular momentum evolution, one can examine the time or redshift evolution and compare formation possibilities.
@@
First we examined the mass–spin diagram in the context of the three black hole mass regimes: stellar-mass black holes formed from gravitational stellar core collapse, intermediate-mass black holes possibly formed through runaway collisions or direct collapse, and supermassive black holes from light and heavy seeds. We approximate redshift evolution using physically motivated mass and spin prescriptions, incorporating both super-Eddington accretion and stochastic processes. These black hole trajectories serve as tracks in the mass–spin space, shaped by their accretion history and seed properties. For instance, stellar-mass BHs can spin up efficiently via fallback or disk accretion, while intermediate-mass black holes undergo diverse accretion and merger episodes, and supermassive black holes experience coherent spin-up in the early Universe and hierarchical galaxy mergers. The mass–spin trajectories carry a fossil record of evolutionary mechanisms, analogous to stellar tracks and clusters in the HR diagram, motivating inference of a black hole ’main sequence’ driven by accretion through cosmic evolution.
@@
Additionally, we provide a zoomed-in view on the stellar-mass black hole regime in a mass–spin diagram constructed with data from the COMPAS binary evolution code. We post-process the black hole data to obtain their spins with three models for the evolution of Wolf-Rayet tidal evolution in progenitor binaries of black-hole mergers. The mass–spin diagram exposes differences in modeling assumptions motivating development of comprehensive and consistent spin evolution models. This will be most useful in the era of future terrestrial GW detectors, where catalogs of black hole binaries are expected to expand by at least 100 times in size [121, 122], consistent with the size of the populations utilized here.
@@
We apply unsupervised and supervised machine learning for classification of a series of realistic and progressively challenging datasets. First we show how simple unsupervised algorithms can recover features in the mass–spin space, such as the intermediate-mass gap with k-means clustering and commonly used mass boundaries in a mass continuum with spectral clustering. Moving beyond simple clustering, we employ deep learning with a variational autoencoder (VAE), a type of neural network, to learn low-dimensional latent embeddings of three synthetic datasets for black hole masses and spins. After training the VAEs on standardized features using PyTorch, we extract the 2D latent spaces and apply an unsupervised algorithm, k-means clustering, and a supervised algorithm, random forest, for machine learning classification. In all three cases, the k-means approach does not learn the latent space representation and fails to accurately cluster the black holes according to their mass–spin correlations. This is clearly shown in Fig. 7 where the unsupervised latent space learning is shown in the top row panels compared to the supervised learning in the bottom panels for true (left column panels) vs predicted (right column panels). The random forest performs best for the first dataset, which is composed of two classes of black hole spins from models with differing assumptions about Wolf-Rayet tidal evolution, with near perfect accuracy and high ARI/NMI. This is because the two classes exhibit more distinct correlations in the mass–spin space compared to the other two datasets which have highly overlapping classes. In the case of dataset 2, this is simply the outcome of comparing raw simulation output from two state-of-the art models of black-hole binary formation channels; in the case of dataset 3, this is by construction to approximate an observed correlation in LVK data.
@@
These results show the utility of the mass–spin diagram for quantifying differences between spin evolution models and distinguishing formation channels of black-hole binaries. Our application of machine learning classification to the diagram reveals interesting challenges, motivating the development of advanced techniques to learn astrophysical datasets. Future work will identify how the motion of black holes across the mass–spin diagram can aid in their classification.